OpenStudy (shelby1290):

I need help with graphing these linear equations: 2x+3y=-12 2x-y=-4

jimthompson5910 (jim_thompson5910):

if x = 0 for 2x+3y=-12, then what is the value of y?

OpenStudy (shelby1290):

2(0)+3y=-12

jimthompson5910 (jim_thompson5910):

that 2(0) turns into 0 when you add 0 to 3y, it stays as 3y

jimthompson5910 (jim_thompson5910):

so we have 3y = -12

OpenStudy (shelby1290):

3y=-12 3 divided by 3 =1 -12 divided by 3 = -4

OpenStudy (shelby1290):

now what do i have to do

OpenStudy (shelby1290):

@jim_thompson5910 ^^

jimthompson5910 (jim_thompson5910):

yes, so if 3y = -12, then y = -4

jimthompson5910 (jim_thompson5910):

together we have x = 0 and y = -4

jimthompson5910 (jim_thompson5910):

so the point (0,-4) lies on 2x+3y=-12

jimthompson5910 (jim_thompson5910):

now solve for x when y = 0

OpenStudy (shelby1290):

2x+3(0)=-12

OpenStudy (shelby1290):

2x=-12

OpenStudy (shelby1290):

2 divided by 2 is 1 and -12 divided by 2 is -6

jimthompson5910 (jim_thompson5910):

yep, so we know that (-6,0) lies on the line 2x+3y=-12

jimthompson5910 (jim_thompson5910):

the two points (0,-4) and (-6,0) lie on 2x+3y=-12

jimthompson5910 (jim_thompson5910):

you just need 2 points to construct a straight line

jimthompson5910 (jim_thompson5910):

(0,-4) is the y-intercept (-6,0) is the x-intercept

OpenStudy (shelby1290):

Alright how about for the second one

jimthompson5910 (jim_thompson5910):

you'll do the same as we did for the first equation

jimthompson5910 (jim_thompson5910):

I'll let you try it out for practice

OpenStudy (shelby1290):

okay so y-int x=0 2(0)-y=-4 -1y=-4 -1y divided by -1 is 1. -4 divided by -1 is 4 y=4 (0,4)

OpenStudy (shelby1290):

am i right?

jimthompson5910 (jim_thompson5910):

(0,4) is on 2x-y=-4, correct

OpenStudy (shelby1290):

x-int y=0 2x-0=-4 2x=-4 2 divided by 2 is 1. -4 divided by 2 is -2 x=-2 (-2,0)

jimthompson5910 (jim_thompson5910):

very good, you have the correct points

OpenStudy (shelby1290):

@jim_thompson5910 thanks for the help!

jimthompson5910 (jim_thompson5910):

you're welcome